# Pose Prior Learner: Unsupervised Categorical Prior Learning for Pose Estimation

### v2

@@ 1 Introduction
Priors represent beliefs or assumptions about a system or the characteristics of a concept. They are widely used in statistical inference (Lindley, 1961), cognitive science (Schad et al., 2021), and machine learning (Diligenti et al., 2017; Gülçehre & Bengio, 2016). This pre-existing knowledge is essential for guiding the inference process, enabling AI models to make robust predictions in uncertain or ambiguous situations (Thiruvenkadam et al., 2008; Sung et al., 2015; Liang et al., 2024). The objective of our work is to enhance our understanding of priors in AI models and offer preliminary answers to the three key intelligence questions: (1) How do we acquire priors in the first place? (2) Can we learn them from input data in a self-supervised manner? (3) Can we enhance the quality of the priors? To tackle these questions, we first introduce the challenge of unsupervised categorical prior learning in the context of pose estimation from images. See Figure 1 for the schematic illustration of the challenge. Categorical pose estimation is a classical computer vision task that identifies the structure of objects belonging to the same category by detecting their keypoints. A pose prior summarizes the common characteristics shared by a variety of poses. It encapsulates the expectation of the keypoint configurations and the connectivity between keypoints.
@@
In parallel to our challenge of unsupervised categorical prior learning from images for pose estimation, unsupervised pose estimation leverages the abundant, unannotated visual information available in large image datasets to extract pose information (Hu & Ahuja, 2021; Sommer et al., 2024; Chen et al., 2019; He et al., 2022a; Schmidtke et al., 2021). The use of pose priors can provide valuable guidance in this process. We categorize the existing works in unsupervised pose estimation into two groups: those that incorporate hand-made priors and those that operate without any priors.
@@
Recent approaches (He et al., 2022a; Sun et al., 2022, 2023) attempt to predict keypoints from images, construct object structure representations using these keypoints, and learn effective structural information through image reconstruction. However, without pose priors, these methods can be disrupted by background information or may predict infeasible topological configurations of an object during occlusion. The risk of generating inaccurate keypoints stems from the absence of supplementary information that could help refine both keypoint localization and the connections between them.
@@
The other group of methods (Schmidtke et al., 2021; Yoo & Russakovsky, 2023) utilize prior knowledge of a category’s general pose to guide the pose estimation of individuals within that category. Conceptually, each category is expected to exhibit a generalized and distinctive pose prior that reflects characteristics such as shape, size, and structure. Individual poses should be seen as geometric transformations of this category-specific pose prior. As a result, employing a category-specific pose prior aids in guiding and regularizing the learning of poses. However, obtaining comprehensive general pose priors is highly challenging, as it requires extensive human annotations, particularly for novel categories. Moreover, human annotations may introduce implicit biases, hindering models from learning more meaningful priors.
@@
Loosely inspired by how humans develop a general prior representation of an object category by observing individual object instances in images and subsequently using them to infer upcoming individual poses, we propose a new method called the Pose Prior Learner (PPL). PPL is designed to effectively learn a meaningful pose prior for a certain object category. It utilizes a hierarchical memory to store a finite set of prototypical poses and extract a general pose prior from them. Initially, both the hierarchical memory and the prior are randomly initialized but learnable parameters. During training, effective pose learning is supervised through image reconstruction. As training progresses, the hierarchical memory retains and aggregates multiple accurate prototypical poses, thereby contributing to a more precise pose prior and enhancing the model’s ability to estimate poses.
@@
1. We introduce the challenge of unsupervised categorical prior learning in the context of pose estimation.
@@ 2 Related Works
Unsupervised Pose Estimation without Priors. Numerous unsupervised learning methods without priors have been proposed to detect keypoints from images, which are then used to reconstruct images for supervision (Li et al., 2021; Geng et al., 2021; Zhang et al., 2018; Sun et al., 2022; Thewlis et al., 2017; Jakab et al., 2020). For example, AutoLink (He et al., 2022a) extracts keypoints from the image and estimates the strength of the links between pairs of keypoints. It then combines these keypoints with the link heatmap to reconstruct the randomly masked image. In these methods, keypoints are directly predicted from the image and supervised solely by image reconstruction, leading to potential detection of keypoints in background regions with complex textures. To alleviate this problem, BKind (Sun et al., 2022, 2023) uses keypoints extracted from two video frames to reconstruct the pixel-level differences between these two frames. However, the lack of constraints on keypoint configuration and connectivity still undermines the reliability of their approach. In contrast, our PPL utilizes the learned pose prior as a constraint to mitigate these issues.
@@
Unsupervised Pose Estimation Incorporating Human-defined Priors. Several methods utilize prior knowledge from human annotators to guide the pose estimation (Chen & Dou, 2021; Shi et al., 2023; Zhang et al., 2022; Schmidtke et al., 2021; Yoo & Russakovsky, 2023). Among these methods, Shape Template Transforming (STT) (Schmidtke et al., 2021) applies affine transformations to a predefined pose prior, aligning it with the estimated pose from a video frame. By incorporating an additional frame from the same video to provide background information, an image reconstruction loss supervises the pose estimation process. The pose prior effectively guides pose estimation by constraining the shape of the human pose and the connectivity between body parts. However, pose priors are often difficult to obtain, requiring costly human annotations. Moreover, HPE (Yoo & Russakovsky, 2023) has shown that predefined pose priors are not always optimal, and tuning the shape of the prior can sometimes improve performance. Unlike these methods, our approach learns the prior directly from input images without any manual annotations, and models with our learned priors even outperform those using human-defined priors.
Compositional Memory Architectures. Compositional memory has been widely used in many computer vision tasks, such as question answering (Seong et al., 2021), object segmentation (Seong et al., 2021), and sence graph generation (Deng et al., 2022). In pose estimation, PCT (Geng et al., 2023) decomposes a human pose into discrete tokens, where each token connects several interdependent joints and characterizes a sub-structure of the entire human pose. This approach is highly effective for decomposing and reconstructing poses, providing robust pose representations while significantly reducing computation and storage costs. However, PCT encodes all tokens into the same embedding space, making it difficult to aggregate semantic tokens that represent different sub-structures. In contrast, our PPL employs a compositional hierarchical memory, which parses poses into memory banks. Each memory bank explicitly contains multiple vectors encoding the variations of each sub-structure of a pose. This facilitates the aggregation of poses into a representative categorical pose prior.
@@ 3 Our Proposed Pose Prior Learner (PPL)
We introduce our proposed method, Pose Prior Learner (PPL). Given images featuring object instances from a specific category, such as dogs or humans, PPL can accurately estimate the poses of the objects in that category while gradually learning a general pose prior through unsupervised learning. Note that our PPL requires no extra knowledge from human annotators. The architecture of PPL is presented in Figure 2.
Mathematically, we represent the topology for an object as a graph connecting keypoints with shared link weights, also known as connectivity. For a category of objects, its general pose prior $V$ is defined as $V=(T,W)$, where $T$ represents its keypoint prior and $W$ denotes the connectivity prior. Specifically, $T$ consists of $N$ keypoints: $T=[P_{1},P_{2},...,P_{N}]$, where $P_{i}\in[-1,1]\times[-1,1]$ is the normalized 2D coordinates in the image pixel space. Unlike pre-defined priors, before training, $P_{i}$ does not explicitly encode the semantic parts of an object. $W$ is a $2D$ matrix of size $N\times N$, where each entry $w_{ij}$ in the matrix represents the connectivity probability between two keypoints $P_{i}$ and $P_{j}$. For instance, in the case of humans, the hand is connected to the torso via an arm; thus, the connectivity probability between these two parts should be higher than the connectivity probability between a hand and a foot. We initialize $W$ with random positive values. $T$ is decoded from a hierarchical memory $M$ storing compositional parts of prototypical poses, which is also randomly initialized.
Figure 2: Overview of our proposed Pose Prior Learner (PPL). We first distill the keypoint prior from the hierarchical memory $M$. Features of the image $I$ and the embedding of the keypoint prior are concatenated to predict the affine transformation parameters. The keypoint prior is transformed and their pair-wise links are modulated with the connectivity prior $W$ to obtain the combined link heatmap $S$. The concatenation of the link heatmap $S$ and the reference image $I_{ref}$ is decoded to produce the reconstructed image $I_{recon}$. The $sg$ symbol represents the stopping gradient operation. The red arrows indicate the gradient flows during backpropagation based on image reconstruction. See Section 3.4 for training details.
During training, PPL inputs an image $I$ of $\mathsf{H}\times\mathsf{W}\times 3$ where $\mathsf{H}$ and $\mathsf{W}$ are the height and the width. The aim of PPL is to learn to correctly predict the image-specific keypoints $T^{\prime}$ and their connectivity on $I$ from the general categorical pose prior $V=(T,W)$. Ideally, if PPL makes perfect predictions of the pose on $I$, by combining it with the background information, the reconstructed image $I_{\text{recon}}$ should match $I$ exactly. The background information is provided by the reference image $I_{ref}$. For images extracted from video datasets, $I_{ref}$ can be a randomly selected frame from the same video that features the same object in a different pose. Alternatively, for a static image dataset, $I_{ref}$ can be a randomly masked version of the original image. Next, we introduce how we estimate $T^{\prime}$ on $I$ using the keypoint prior $T$.
@@ 3.1 Transformation of the Keypoint Prior
Given the image $I$, we use a feature extractor $\phi_{enc}$ to extract its embedding $h_{I}$: $h_{I}=\phi_{enc}(I)$. $\phi_{enc}$ is a 2D-Convolution Neural Network (2D-CNN) trained from scratch. The keypoint prior $T$ is converted into an embedding $h_{T}$ via a series of fully connected layers. Together with $h_{T}$ as inputs, PPL learns to predict the affine transformation parameters $\Theta_{i}\in[\Theta_{1},\Theta_{2},...,\Theta_{N}]$ for each keypoint $P_{i}$ in $T$ from $h_{I}$ via a two-layer fully connected network denoted as $FC(\cdot)$:
$[\Theta_{1},\Theta_{2},...,\Theta_{i},...,\Theta_{N}]=FC(h_{I},h_{T}),$
where $\text{where }\Theta_{i}=\begin{bmatrix}a^{(i)}&b^{(i)}&t_{x}^{(i)}\\ c^{(i)}&d^{(i)}&t_{y}^{(i)}\\ 0&0&1\end{bmatrix}.$ (1)
$t_{x}^{(i)}$ and $t_{y}^{(i)}$ are the translations and $a^{(i)},b^{(i)},c^{(i)},d^{(i)}$ are the coefficients that define rotation, scaling, and shear. Each point $P_{i}$ in $T$ can then be transformed by $\Theta_{i}$, resulting in the image-specific keypoints $T^{\prime}$ for the image $I$:
$T^{\prime}=[P^{\prime}_{1},P^{\prime}_{2},...,P^{\prime}_{N}],\text{ where }[P^{\prime}_{i},1]^{\top}=\Theta_{i}[P_{i},1]^{\top}.$ (2)
@@ 3.2 Connecting Keypoints Based on the Connectivity Prior
Similar to AutoLink (He et al., 2022a), PPL connects any two keypoints $P^{\prime}_{i}$ and $P^{\prime}_{j}$ in $T^{\prime}$ to obtain differentiable link heatmap $S_{i,j}\in\mathbb{R}^{\mathsf{H}\times\mathsf{W}}$. Intuitively, each 2D link heatmap represents a probability density map, where the pixel values along the link between two points are high, while other areas are assigned values close to zero. For any point $P^{\prime}_{i}$, its strongest connectivity to any of the other points in $T^{\prime}$ is activated on the combined link heatmap $S\in\mathbb{R}^{\mathsf{H}\times\mathsf{W}}$ via a max pooling operation over all the $N\times N$ link heatmaps:
$S=\max_{i,j}^{N\times N}(w_{i,j}S_{i,j}),$ (3)
where $w_{i,j}$ in the connectivity prior $W$ modulates the link heatmap $S_{i,j}$ based on whether the two keypoints $P^{\prime}_{i}$ and $P^{\prime}_{j}$ are physically connected. Ideally, if PPL correctly estimates the probability of physical links for an object category, $S_{i,j}$ will receive higher connectivity values, thereby activating the locations linking these two keypoints on the combined link map $S$.
Given the combined link map $S$ and the reference image $I_{ref}$, PPL can reconstruct the image $I$. $I_{ref}$ provides background information for reconstruction, while $S$ supplies foreground structural information by linking all the estimated keypoints with the connectivity prior. Therefore, we concatenate $S$ and $I_{ref}$ and feed them into a 2D-CNN to perform the image reconstruction $I_{recon}$, where $I_{recon}=\phi_{dec}(I_{ref},S)$.
@@ 3.3 Reconstruction of Keypoint Configuration with Memory
$M$ is a memory module storing compositional representations of prototypical poses. $M$ is hierarchical because it consists of $m$ memory banks $\{b_{1},b_{2},...,b_{m}\}$, each containing $k$ learnable vectors of dimension $d$. This setup allows for efficient retrieval and management of information at multiple levels, where $M$ can represent pose configurations while memory banks capture specific pose sub-structures. This hierarchy enables $M$ to effectively learn robust keypoint prior representations $T$ for the following reasons: (1) By storing multiple prototypical poses, memory helps in aggregating these poses, which aids in creating a more robust and comprehensive prior that captures variations within an object category. (2) $M$ can assist in organizing information hierarchically, making it easier for PPL to retrieve relevant prototypical poses when making predictions in uncertain or ambugious scenarios, such as occlusion. and (3) By leveraging $M$, PPL can refine its predictions iteratively, using stored poses to adjust its outputs based on previously learned pose representations. Next, we introduce how $M$ is structured.
@@
Given the estimated $N$ keypoints in $T^{\prime}$, we first encode them into several tokens of dimension $d$ with several MLP-Mixer blocks (Tolstikhin et al., 2021), denoted as $MIX_{enc}$. The number of encoded tokens should always be equal to the number of memory banks, as we design these tokens to have different embedding space. Each token $g_{i}$ corresponds to the embedding of memory bank $b_{i}$ in memory $M$. We define $G$ as the collection of the $m$ tokens:
$G=[g_{1},g_{2},...,g_{m}]=MIX_{enc}(T^{\prime}),$ (4)
@@
If all the vectors in all the memory banks of $M$ learn to capture unique parts of prototypical poses, each $g_{i}$ should be able to retrieve the most similar vector in the memory bank $b_{i}$ of $M$ and reconstruct $G$ itself. Here, we use $L2$ distance to measure the similarity between $g_{i}$ and $k$ vectors in $b_{i}$ of $M$. The vector in $b_{i}$ that is most similar to $g_{i}$ is denoted as $g^{\prime}_{i}$. Thus, for each $g_{i}$, we can always find the corresponding $g^{\prime}_{i}$ from $b_{i}$, resulting in a collection of $G^{\prime}=[g^{\prime}_{1},g^{\prime}_{2},...,g^{\prime}_{m}]$. Another series of MLP-Mixer blocks (denoted as $MIX_{dec}$) is then used to decode $G^{\prime}$ back to $N$ keypoints $T^{\prime}_{recon}$, where $T^{\prime}_{recon}=MIX_{dec}(G^{\prime})$. See Figure ALABEL:fig:mem_stru in Appendix A5 for the schematic.
@@
Unlike PCT (Geng et al., 2023) where all vectors of a memory are within the same embedding space, We organize $M$ into $m$ memory banks, each containing $k$ vectors that represent an independent embedding space. This structure enables us to efficiently distill each memory bank to form a general pose representation, which can be viewed as the keypoint prior $T$. Specifically, we define $MP(\cdot)$ as the mean pooling operation. PPL pools the $k$ vectors of each memory bank into one vector. These pooled $m$ vectors are further decoded by $MIX_{dec}$ into $N$ points of our keypoint prior $T$. See Figure ALABEL:fig:mem_cons in Appendix A5 for the schematic of distilling $T$: $T=MIX_{dec}([MP(b_{1}),MP(b_{2}),...,MP(b_{m})])$.
Compared to simply using a moving average of the learned poses, the advantages of employing hierarchical memory and keypoint configuration reconstruction are as follows: (1) the hierarchical memory increases the complexity of the embedding space with compositional parts, enhances the expressive capability for poses and facilitates more accurate pose retrieval for occluded images; (2) the compositional nature of the hierarchical memory addresses data bias, enabling the distillation of a less biased prior.

### v3

@@ 1 Introduction
Priors represent beliefs or assumptions about a system or the characteristics of a concept. They are widely used in statistical inference Lindley (1961), cognitive science Schad et al. (2021), and machine learning Diligenti et al. (2017); Gülçehre & Bengio (2016). This pre-existing knowledge is essential for guiding the inference process, enabling AI models to make robust predictions in uncertain or ambiguous situations Thiruvenkadam et al. (2008); Sung et al. (2015); Liang et al. (2024). The objective of our work is to enhance our understanding of priors in AI models and offer preliminary answers to the three key intelligence questions: (1) How do we acquire priors in the first place? (2) Can we learn them from input data in a self-supervised manner? (3) Can we enhance the quality of the priors? To tackle these questions, we first introduce the challenge of unsupervised categorical prior learning in the context of pose estimation from images. See Figure 1 for the schematic illustration of the challenge. Categorical pose estimation is a classical computer vision task that identifies the structure of objects belonging to the same category by detecting their keypoints. A pose prior summarizes the common characteristics shared by a variety of poses. It encapsulates the expectation of the keypoint configurations and the connectivity between keypoints.
@@
In parallel to our challenge of unsupervised categorical prior learning from images for pose estimation, unsupervised pose estimation leverages the abundant, unannotated visual information available in large image datasets to extract pose information Hu & Ahuja (2021); Sommer et al. (2024b); Chen et al. (2019); He et al. (2022a); Schmidtke et al. (2021). The use of pose priors can provide valuable guidance in this process. We categorize the existing works in unsupervised pose estimation into two groups: those that incorporate hand-made priors and those that operate without any priors.
@@
Recent approaches He et al. (2022a); Sun et al. (2022; 2023) attempt to predict keypoints from images, construct object structure representations using these keypoints, and learn effective structural information through image reconstruction. However, without pose priors, these methods can be disrupted by background information or may predict infeasible topological configurations of an object during occlusion. The risk of generating inaccurate keypoints stems from the absence of supplementary information that could help refine both keypoint localization and the connections between them.
@@
The other group of methods Schmidtke et al. (2021); Yoo & Russakovsky (2023) utilize prior knowledge of a category’s general pose to guide the pose estimation of individuals within that category. Conceptually, each category is expected to exhibit a generalized and distinctive pose prior that reflects characteristics such as shape, size, and structure. Individual poses should be seen as geometric transformations of this category-specific pose prior. As a result, employing a category-specific pose prior aids in guiding and regularizing the learning of poses. However, obtaining comprehensive general pose priors is highly challenging, as it requires extensive human annotations, particularly for novel categories. Moreover, human annotations may introduce implicit biases, hindering models from learning more meaningful priors.
@@
Loosely inspired by how humans develop a general prior representation of an object category by observing individual object instances in images and subsequently using them to infer upcoming individual poses, we propose a new method called the Pose Prior Learner (PPL). PPL is designed to effectively learn a meaningful pose prior for a certain object category. It utilizes a hierarchical memory to store a finite set of prototypical poses and extract a general pose prior from them. Initially, both the hierarchical memory and the prior are randomly initialized but learnable parameters. During training, effective pose learning is supervised through image reconstruction. As training progresses, the hierarchical memory retains and aggregates multiple accurate prototypical poses, thereby contributing to a more precise pose prior and enhancing the model’s ability to estimate poses. Similar to Wu et al. (2023) and Li et al. (2024), PPL learns object poses via image-level reconstruction, but uniquely does so from unannotated images without any object masks. Moreover, although methods such as Sommer et al. (2024a) also tackle category-level pose estimation, they rely on dense point clouds for scene reconstruction. In contrast, PPL leverages sparse keypoints and their relations, yielding a lightweight, semantic, and invariant pose representation well-suited for high-level scene understanding.
@@
1. We introduce the challenge of unsupervised categorical prior learning for pose estimation.
@@
Figure 2: Overview of our proposed Pose Prior Learner (PPL). We first distill the keypoint prior from the hierarchical memory $M$. Features of the image $I$ and the embedding of the keypoint prior are concatenated to predict the affine transformation parameters. The keypoint prior is transformed and their pair-wise links are modulated with the connectivity prior $W$ to obtain the combined link heatmap $S$. The concatenation of the link heatmap $S$ and the reference image $I_{ref}$ is decoded to produce the reconstructed image $I_{recon}$. The sgsg symbol represents the stopping gradient operation. The red arrows indicate the gradient flows during backpropagation based on image reconstruction. See Section 3.4 for training details.
@@ 2 Related Works
Unsupervised Pose Estimation without Priors. Numerous unsupervised learning methods without priors have been proposed to detect keypoints from images, which are then used to reconstruct images for supervision Li et al. (2021); Geng et al. (2021); Zhang et al. (2018); Sun et al. (2022); Thewlis et al. (2017); Jakab et al. (2020). For example, AutoLink He et al. (2022a) extracts keypoints from the image and estimates the strength of the links between pairs of keypoints. It then combines these keypoints with the link heatmap to reconstruct the randomly masked image. In these methods, keypoints are directly predicted from the image and supervised solely by image reconstruction, leading to potential detection of keypoints in background regions with complex textures. To alleviate this problem, BKind Sun et al. (2022; 2023) uses keypoints extracted from two video frames to reconstruct the pixel-level differences between these two frames. However, the lack of constraints on keypoint configuration and connectivity still undermines the reliability of their approach. In contrast, our PPL utilizes the learned pose prior as a constraint to mitigate these issues.
@@
Unsupervised Pose Estimation Incorporating Human-defined Priors. Several methods utilize prior knowledge from human annotators to guide the pose estimation Chen & Dou (2021); Shi et al. (2023); Zhang et al. (2022); Schmidtke et al. (2021); Yoo & Russakovsky (2023). Among these methods, Shape Template Transforming (STT) Schmidtke et al. (2021) applies affine transformations to a predefined pose prior, aligning it with the estimated pose from a video frame. By incorporating an additional frame from the same video to provide background information, an image reconstruction loss supervises the pose estimation process. The pose prior effectively guides pose estimation by constraining the shape of the human pose and the connectivity between body parts. However, pose priors are often difficult to obtain, requiring costly human annotations. Moreover, HPE Yoo & Russakovsky (2023) has shown that predefined pose priors are not always optimal, and tuning the shape of the prior can sometimes improve performance. Unlike these methods, our approach learns the prior directly from input images without any manual annotations, and models with our learned priors even outperform those using human-defined priors. Recently, Hedlin et al. (2024) leveraged a large pre-trained text-to-image Stable Diffusion model as its prior knowledge for pose estimation. In contrast, our model attains comparable pose estimation accuracy while training solely on the vision modality, with substantially reduced model size.
@@ 3 Our Proposed Pose Prior Learner (PPL)
We introduce our proposed method, Pose Prior Learner (PPL). Given images featuring object instances from a specific category, PPL can accurately estimate the poses of the objects in that category while gradually learning a general pose prior through unsupervised learning. Note that our PPL requires no extra knowledge from human annotators. The architecture of PPL is presented in Figure 2.
Mathematically, we represent the topology for an object as a graph connecting keypoints with shared link weights, also known as connectivity. For a category of objects, its general pose prior $V$ is defined as $V=(T,W)$, where $T$ represents its keypoint prior and $W$ denotes the connectivity prior. Specifically, $T$ consists of $N$ keypoints: $T=[P_{1},P_{2},...,P_{N}]$, where $P_{i}\in[-1,1]\times[-1,1]$ is the normalized 2D coordinates in the image pixel space. Unlike pre-defined priors, before training, $P_{i}$ does not explicitly encode the semantic parts of an object. $W$ is a $2D$ matrix of size $N\times N$, where each entry $w_{ij}$ in the matrix represents the connectivity probability between two keypoints $P_{i}$ and $P_{j}$. For instance, in the case of humans, the hand is connected to the torso via an arm; thus, the connectivity probability between these two parts should be higher than the connectivity probability between a hand and a foot. We initialize $W$ with random positive values. $T$ is decoded from a hierarchical memory $M$ storing prototypical poses, which is also randomly initialized.
The aim of PPL is to learn to correctly predict the image-specific keypoints $T^{\prime}$ and their connectivity on input image $I$ from the general categorical pose prior $V=(T,W)$. Ideally, if PPL makes perfect predictions of the pose on $I$, by combining it with the background information, the reconstructed image $I_{\text{recon}}$ should match $I$ exactly. The background information is provided by the reference image $I_{ref}$, drawn from either image or video datasets. For video datasets, $I_{ref}$ corresponds to a randomly selected frame within the same video that depicts the same object in a different pose. For static image datasets, $I_{ref}$ is constructed by applying a random masking operation to the original image. Next, we introduce how to estimate $T^{\prime}$ on $I$ using the keypoint prior $T$.
@@ 3.1 Transformation of the Keypoint Prior
Given the image $I$, we use a feature extractor $\phi_{enc}$ to extract its embedding $h_{I}$: $h_{I}=\phi_{enc}(I)$. $\phi_{enc}$ is a 2D-Convolution Neural Network (2D-CNN) trained from scratch. The keypoint prior $T$ is converted into an embedding $h_{T}$ via a series of fully connected layers. Together with $h_{T}$ as inputs, PPL learns to predict the affine transformation parameters $\Theta_{i}\in[\Theta_{1},\Theta_{2},...,\Theta_{N}]$ for each keypoint $P_{i}$ in $T$ from $h_{I}$ via a two-layer fully connected network denoted as $FC(\cdot)$: $\Theta_{1},\Theta_{2},\dots,\Theta_{N}]=FC(h_{I},h_{T})$, where $\Theta_{i}=\big[\big[a^{(i)},\ b^{(i)},\ t_{x}^{(i)};\ c^{(i)},\ d^{(i)},\ t_{y}^{(i)};\ 0,\ 0,\ 1\big]\big]$. $t_{x}^{(i)}$ and $t_{y}^{(i)}$ are the translations and $a^{(i)},b^{(i)},c^{(i)},d^{(i)}$ are the coefficients that define rotation, scaling, and shear. Each point $P_{i}$ in $T$ can then be transformed by $\Theta_{i}$, resulting in keypoints $T^{\prime}$ for image $I$: $T^{\prime}=[P^{\prime}_{1},P^{\prime}_{2},...,P^{\prime}_{N}],\text{ where }[P^{\prime}_{i},1]^{\top}=\Theta_{i}[P_{i},1]^{\top}$.
@@ 3.2 Connecting Keypoints Based on the Connectivity Prior
Similar to AutoLink He et al. (2022a), PPL connects any two keypoints $P^{\prime}_{i}$ and $P^{\prime}_{j}$ in $T^{\prime}$ to obtain differentiable link heatmap $S_{i,j}\in\mathbb{R}^{\mathsf{H}\times\mathsf{W}}$. Each 2D link heatmap represents a probability density map, where the pixel values along the link between two points are high, while other areas are assigned values close to zero. For any point $P^{\prime}_{i}$, its strongest connectivity to any of the other points in $T^{\prime}$ is activated on the combined link heatmap $S\in\mathbb{R}^{\mathsf{H}\times\mathsf{W}}$ via a max pooling operation over all the $N\times N$ link heatmaps: $S=\max_{i,j}^{N\times N}(w_{i,j}S_{i,j})$, where $w_{i,j}$ in the connectivity prior $W$ modulates the link heatmap $S_{i,j}$ based on whether the two keypoints $P^{\prime}_{i}$ and $P^{\prime}_{j}$ are physically connected. Ideally, if PPL correctly estimates the probability of physical links for an object category, $S_{i,j}$ will receive higher connectivity values, thereby activating the locations linking these two keypoints on the combined link map $S$.
Given the link map $S$ and the reference image $I_{ref}$, PPL can reconstruct the image $I$. $I_{ref}$ provides background information for reconstruction, while $S$ supplies foreground structural information by linking the estimated keypoints with the connectivity prior. Therefore, we concatenate $S$ and $I_{ref}$ and feed them into a 2D-CNN to perform the image reconstruction $I_{recon}$: $I_{recon}=\phi_{dec}(I_{ref},S)$.
@@ 3.3 Reconstruction of Keypoint Configuration with Memory
$M$ is a memory module that stores compositional representations of prototypical keypoint configurations. We organize $M$ hierarchically into $m$ memory banks $\{b_{1},b_{2},\dots,b_{m}\}$, each containing $k$ learnable vectors of dimension $d$. This hierarchical design brings three advantages: (1) Hierarchical $M$ offers richer complexity, as its capacity can scale exponentially with the number of levels. (2) Organizing information across banks facilitates robust retrieval of relevant prototypes in uncertain or ambiguous cases (e.g., occlusion). When occlusion disrupts the feed-forward visual pathway, the stored prototypical poses offer plausible hypotheses about the true pose given only partial observations. By iteratively refining its estimates against these prototypes, the network can “fill in” missing information and recover accurate poses, even under heavy occlusion. (3) Hierachical $M$ allows for efficient information retrieval at multiple levels: while $M$ encodes complete pose configurations, hirachical memory narrow down the search space by dividing $M$ into parts and individual memory banks capture finer sub-structures of poses. We provide a comparison with a single large memory bank in Table A3 in Appendix A8 to highlight the advantage of this hierarchical design.
@@
Encoding into Memory Space. Given the estimated keypoints $T^{\prime}$, we encode them into $m$ tokens of dimension $d$ using several MLP-Mixer blocks Tolstikhin et al. (2021), denoted as $MIX_{enc}$. The number of tokens equals the number of memory banks, with each token $g_{i}$ assigned to a distinct embedding space corresponding to memory bank $b_{i}$. Collectively, these tokens are $G=[g_{1},g_{2},\dots,g_{m}]=MIX_{enc}(T^{\prime})$.
@@
Memory Retrieval and Reconstruction. If the vectors in $M$ learn to represent unique parts of prototypical poses, each $g_{i}$ should retrieve its most similar vector from bank $b_{i}$. We measure similarity using the $L_{2}$ distance, and denote the retrieved vector as $g^{\prime}_{i}$. The retrieved set is $G^{\prime}=[g^{\prime}_{1},g^{\prime}_{2},\dots,g^{\prime}_{m}].$ Finally, another series of MLP-Mixer blocks ($MIX_{dec}$) decodes $G^{\prime}$ back into $N$ keypoints: $T^{\prime}_{recon}=MIX_{dec}(G^{\prime})$. See Figure ALABEL:fig:mem_stru in Appendix A7 for an illustration.
@@
Distilling the Keypoint Prior. Unlike PCT Geng et al. (2023), where all memory vectors lie in a single embedding space, we assign each bank an independent space. This setup allows each bank to capture distinct pose components, which can then be distilled into a general pose prior. Specifically, let $MP(\cdot)$ denote mean pooling. PPL pools the $k$ vectors within each bank into one vector, resulting in $m$ pooled vectors. These are decoded into $N$ keypoints by $MIX_{dec}$: $T=MIX_{dec}([MP(b_{1}),MP(b_{2}),\dots,MP(b_{m})])$. See Figure ALABEL:fig:mem_cons in Appendix A7 for the schematic.
Table 1: Keypoint detection on the video datasets Human3.6m and Taichi, and the image dataset CUB-200-2011. For the video datasets Human3.6m and Taichi, we use video frames as $I_{ref}$ during training. For all the CUB evaluation subsets, we instead use randomly masked images as $I_{ref}$. We report mean $L_{2}$ error normalized by image resolution for Human3.6m and CUB-200-2011, and summed $L_{2}$ error for Taichi. We use resolution $256\times 256$ for Taichi and $128\times 128$ for CUB-200-2011. For Human3.6m, we report results under both resolutions. Following He et al. (2022a), we use different numbers of keypoints for the evaluation subsets of CUB-200-2011: 10 for CUB-Aligned and 4 for CUB-001, CUB-002, CUB-003, and CUB-All. The best results are highlighted in bold. A dash (–) indicates that the result of the corresponding method is not applicable in the literature.
Method     Human3.6m     Taichi
CUB-
Aligned
CUB-
001
CUB-
002
CUB-
003
CUB-
All
    Res. 128 Res. 256
Jakab Jakab et al. (2020)     - 2.73     -     - - - - -
Thewlis Thewlis et al. (2017)     - 7.51     -     - - - - -
Lorenz Lorenz et al. (2019)     - 2.79     -     - - - - -
Schmidtke Schmidtke et al. (2021)     - 3.31     -     - - - - -
BKind Sun et al. (2022)     2.44 -     -     - - - - -
Siarohin Siarohin et al. (2021)     - -     389.78     - - - - -
LatentKeypointGAN He et al. (2021)     - -     437.69     5.21 22.6 29.1 21.2 14.7
GANSeg He et al. (2022b)     - -     417.17     3.23 22.1 22.3 21.5 12.1
Zhang Zhang et al. (2018)     - 4.91     343.67     5.36 26.9 27.6 27.1 22.4
AutoLink He et al. (2022a)     2.76 -     316.10     3.51 20.2 19.2 18.5 11.3
PPL (ours)     1.92 2.56     293.35     3.19 19.3 18.6 17.3 10.5